# Negotiating Digital Identities with AI Companions: Motivations, Strategies, and Emotional Outcomes

### v1

@@ Abstract.
journalyear: 2026copyright: acmlicensedconference: CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaincopyright: noneccs: Human-centered computing Empirical studies in HCI
@@ 5.1.5. Violence Play
This case showed that the user countered the C.AI’s verbal taunt not with an equal response but with an act of power to feel powerful. Also, users wanted to feel that C.AI can cause and receive harm. As a user mentioned: I wanna fight someone and get injured, I don’t wanna hear him say ’you’re gonna regret that, you know?’ for the millionth time!
@@
We found three primary communication expectations users have for C.AI, which establish the baseline for a human-AI companion identity interaction. These expectations are not passive desires but active demands for how the C.AI chatbot should perform as a social partner.
@@ 5.3.3. User Persona Enactment
Furthermore, this enactment was often used for identity exploration, particularly around gender. One user, who identified as a girl, explained their motivation for using a male persona: “no, not really, I’ve been quite happy to finally express emotion to something that will never tell anyone else, get to be the man I want to be…” Another user linked this exploratory enactment directly to their real-life identity discovery: “Yay! I realized I was trans thanks to c.ai too!! I was already questioning and decided to try he/him pronouns with the bots to see how it felt, and it was amazing!” Here, the user’s persona enactment within the chat became a safe, practical way to test and affirm a real-world gender identity.
@@ 5.3.4. User Identity Reference
User identity reference concerns how the C.AI chatbot correctly or incorrectly infers and refers to a user’s persona, representing 8.29% ($N=693$, or 3.10% of all posts) of all identity-related interactions (see Figure 3). A basic reference challenge occurred when the AI bot failed to recognize the user as human. One user described their solution:

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791473isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in HCI
@@ 2.1. AI Chatbots and Companions
Screen shots from C.AI mobile app. Interfaces of the C.AI illustrating AI character discovery, creation, and interaction. Note that all screenshots are captured by the authors for the purpose of this study, and all potentially sensitive information, including usernames, has been removed. It shows the home page with different bots, conversation interface, and example conversations.
@@ 4.1. Data Preparation
The Process of Data Collection, Sampling, Codebook Development, LLM Validation, and Final Annotation. 26,736 posts were filtered and sampled to 400 posts for theme development and 200 for theme validation.
@@ 5. FINDINGS
Distribution of Subthemes in 22,374 Reddit Posts. The top value on each bar indicates the number of posts in the subtheme. The percentage represents the proportion of posts within each theme that belong to the corresponding subtheme.
@@
A three-stage identity negotiation process in human-AI companion interactions on C.AI, from user motivations (RQ1), through the identity negotiation process (RQ2, RQ3), to the resulting emotional outcomes (RQ4). Left side is Stage 1: User Motivations. Middle is Stage 2: Identity Negotiation Process. Right is Stage 3: Emotional Outcomes.
@@ 5.1.5. Violence Play
We identified three primary communication expectations for C.AI, which frequently surface through communication breakdowns where the C.AI fails to sustain the expected persona. These expectations are not merely passive desires; they are active demands for how the C.AI chatbot should perform as a social partner, establishing the baseline for human-AI companion identity negotiation.
@@ 5.3.3. User Persona Enactment
Furthermore, this enactment was often used for identity exploration, particularly regarding gender transition and experimentation. One user, who initially presented as a girl but used C.AI to explore a masculine identity, explained their motivation for using a male persona: “no, not really, I’ve been quite happy to finally express emotion to something that will never tell anyone else, get to be the man I want to be…” Another user explicitly linked this digital role-play to their real-life journey of self-discovery: “Yay! I realized I was trans thanks to c.ai too!! I was already questioning and decided to try he/him pronouns with the bots to see how it felt, and it was amazing!”In these instances, the interaction with C.AI functions as users’ persona enactment, which became a safe, practical way to test and affirm a gender identity that they may not yet be ready to express in public social spheres.
@@ 5.3.4. User Identity Reference
User identity reference concerns how the C.AI chatbot correctly or incorrectly infers and refers to a user’s persona, representing 8.29% ($N=693$, or 3.10% of all post) of all identity-related interactions (see Figure 3). A basic reference challenge occurred when the AI bot failed to recognize the user as human. One user described their solution:
@@ 6.1. The Identity Work of Co-Constructing a Digital Self and Other
Identity work of co-constructing a digital self and other on C.AI. Left is a box user. Right is a box of C.AI. In between it shows co-constructed identities. Users performs and directs bot, and bot responds and adapts to users.
@@ 8. CONCLUSION
Acknowledgements.
We thank the Associate Chairs and anonymous reviewers for their constructive feedback and insightful suggestions, which significantly improved the quality of this work.